Two plates are placed at a distance of 0.15 mm apart. The lower plate is fixed while the upper plate having surface area 1.0 m^2 is pulled at 0.3 m/s. Find the force and power required to maintain this speed, if the fluid separating them is having viscosity 1.5 poise.
1
Expert's answer
2020-11-09T20:28:22-0500

Let force required be "F"


"F=\nu\cdot(area)\cdot\dfrac{dv}{dz}"

Putting all the values in SI unit


"F=1.5\times(0.1Pa\cdot s)\cdot(1.0m^2)\cdot\dfrac{0.3\times10^{-3}m/s}{0.15\times10^{-3}m}="

"=0.3N"

The force is "0.3N."


Let power required to maintain this speed be "P"


"P=F\cdot v"

"P=0.3N\cdot0.3\times10^{-3}m/s=9\times10^{-5}W=90\mu W"

Power required to maintain this speed is "9\times10^{-5}W=90\mu W."